The difference of the areas of two squares is 75 square feet. Each side of the larger square is twice the length of the smaller square. Find the length of a side of each square.: The difference of the areas of two squares is 75 square feet. Each side of the larger square is twice the length of the smaller square. Find the length of a side of each square.
Answer by ankor@dixie-net.com(4675)

The difference of the areas of two squares is 75 square feet. Each side of the
larger square is twice the length of the smaller square. Find the length of a
side of each square.
:
Let x = length of the side of the smaller square
then
2x = length of the side of the larger square
:
:
Large sq area - small sq area = 75 sq/ft
(2x)^2 - x^2 = 75
:
4x^2 - x^2 = 75
:
3x^2 = 75
x^2 = 75/3
x^2 = 25
x = 5 ft side of the small square
and
10 ft side of the large square
:
Check solution
10^2 - 5^2 = 75